40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 21 12:25:46 2019, gautam, Update, ALS, DFD electronics checkout ALSnoiseAnomaly.pdfDFDcalib.pdfspectra.pdfPTgains.pdf
    Reply  Tue Oct 22 15:32:15 2019, gautam, Update, ALS, EX uPDH electronics checkout uPDH_X_OLTFs.pdf
       Reply  Wed Oct 23 10:23:26 2019, rana, Update, ALS, EX uPDH electronics checkout 
          Reply  Wed Oct 23 11:11:01 2019, gautam, Update, ALS, EX uPDH electronics checkout 
    Reply  Fri Oct 25 01:04:49 2019, gautam, Update, ALS, ALS electronics chain was saturating 
       Reply  Mon Oct 28 23:20:11 2019, gautam, Update, ALS, ALS power budget 
Message ID: 14984     Entry time: Tue Oct 22 15:32:15 2019     In reply to: 14981     Reply to this: 14986
Author: gautam 
Type: Update 
Category: ALS 
Subject: EX uPDH electronics checkout 

Summary:

The EX PDH setup had what I thought was insufficient phase and gain margins. So I lowered the gain a little - the price paid was that the suppression of laser frequency noise of the end laser was reduced. I actually think an intermediate gain setting (G=7) can give us ~35 degrees of phase margin, ~10dB gain margin, and lower residual unsuppressed AUX laser noise - to be confirmed by measurement later. See here for the last activity I did - how did the gain get increased? I can't find anything in the elog.

Attachment 1: uPDH_X_OLTFs.pdf  174 kB  Uploaded Tue Oct 22 18:22:27 2019  | Hide | Hide all
uPDH_X_OLTFs.pdf
ELOG V3.1.3-